OpenStudy (anonymous):

HELP !! .. GIVING OUT MEDALS :) Evaluate 9x + 3xy - 2 for x = 3 and y = -3

OpenStudy (anonymous):

(9*3)+(3*3*-3)-2 = -2

OpenStudy (anonymous):

\[9(3)+3(3\times-3) -2= -2\]